OpenStudy (xoxolove456):

Sam can mow a lawn one hour faster than ally can. Together they can mow the lawn in 2.5 hours. How long would it take each person working alone??

OpenStudy (dumbcow):

\[\frac{1}{s} +\frac {1}{a} = \frac{1}{2.5}\] \[s = a-1\] substitute and solve for "a"